# Effortless, Simulation-Efficient Bayesian Inference using Tabular Foundation Models

### v1

@@ D.6 Order of autoregressive sampling
As in Sec. 3.1, we compute the C2ST with respect to the ground truth for evaluation. We run NPE-PF for the default autoregressive order (as defined in the SBI benchmark or in the equations above) and up to ten (random) permutations over three random seeds. Note that tasks with two or three dimensions have only two or six possible permutations, respectively (including the default one), so random subsampling is not required.
@@
Figure D-7: Autoregressive ordering. C2ST across varying simulation budgets for two additional synthetic tasks, and five tasks from the SBI benchmark that are not permutation invariant. Blue lines indicate the average C2ST accuracy over three seeds for up to ten (random) permutations of the sampling order. Dashed orange lines indicate the default order.
@@
Across all benchmark tasks and simulation budgets, NPE-PF achieves nearly identical performance across the default and permuted orders (Fig. D-7). That is, for these benchmark tasks, the performance of NPE-PF is not affected by permuting dimensions. These results show that NPE-PF is not sensitive to the order in which the parameters are sampled. As a result, users need not worry about providing NPE-PF with an “optimal” order. While differences in performance may be possible for very high dimensions or artificial examples, for the applications we consider here, an arbitrary permutation is sufficient.
@@
Figure D-8: Posterior predictives of TSNPE-PF for real observations from the Allen cell type database and a simulation budget of $10^{4}$. Note that the inference is not performed directly on the action potential time series but on seven summary statistics computed from it.
@@
Figure D-9: Posterior predictives of the TSNPE baseline for real observations from the Allen cell type database and a simulation budget of $10^{4}$. Note that the inference is not performed directly on the action potential time series but on seven summary statistics computed from it.
@@
Figure D-10: Pyloric simulator. Posterior distributions for all 31 parameters estimated by TSNPE-PF.

### v2

@@ D.8 Order of autoregressive sampling
As in Sec. 3.1, we compute the C2ST with respect to the ground truth for evaluation. We run NPE-PFN for the default autoregressive order (as defined in the SBI benchmark or in the equations above) and up to ten (random) permutations over three random seeds. Note that tasks with two or three dimensions have only two or six possible permutations, respectively (including the default one), so random subsampling is not required.
@@
Figure D-9: Autoregressive ordering. C2ST across varying simulation budgets for two additional synthetic tasks, and five tasks from the SBI benchmark that are not permutation invariant. Blue lines indicate the average C2ST accuracy over three seeds for up to ten (random) permutations of the sampling order. Dashed orange lines indicate the default order.
@@
Across all benchmark tasks and simulation budgets, NPE-PFN achieves nearly identical performance across the default and permuted orders (Fig. D-9). That is, for these benchmark tasks, the performance of NPE-PFN is not affected by permuting dimensions. These results show that NPE-PFN is not sensitive to the order in which the parameters are sampled. As a result, users need not worry about providing NPE-PFN with an “optimal” order. While differences in performance may be possible for very high dimensions or artificial examples, for the applications we consider here, an arbitrary permutation is sufficient.
D.9 Unconditional density estimation on the UCI datasets
Here, we apply TabPFN on some classical unconditional density estimation benchmark tasks from the UCI repository [56] as used in several other works [41, 42, 97]. Note that while we still perform density estimation here, it does not involve posterior distributions. Thus, we do not refer to this approach as NPE-PFN. Specifically, we consider the Gas, Power, Hepmass, and Miniboone datasets. These tabular datasets range in dimensionality from 6 to 43 features and contain between $3.1\cdot 10^{4}$ and over 1 million samples. We here investigate the unconditional density estimation performance of TabPFN in the low sample regime in comparison to a neural spline flow (NSF) [42]. As in previous works, we compute the negative log-likelihood (NLL) under a held-out test set. While we evaluate on the full test sets, we only use $10^{3}$, $10^{4}$, or (if applicable) $10^{5}$ samples for training. For the two larger settings, we use partitioning (Sec. B.4) with 10 clusters when estimating densities with TabPFN.
Figure D-10: Results for unconditional density estimation on the UCI datasets. Negative-log-likelihood (NLL) for TabPFN (with 1 and 10 clusters) and the neural spline flow (NSF) across the different UCI datasets. Dots indicate averages, and bars show standard deviation over five independent runs.
For $10^{2}$ and $10^{3}$ samples, TabPFN achieves a smaller negative log-likelihood on all four datasets (Fig. D-10). Similarly, for the $10^{4}$ samples, TabPFN outperforms NSF on all but the power dataset. Interestingly, for the lower dimensional datasets, clustering further reduces the NLL despite having access to the same total number of samples. On the other hand, for the higher dimensional datasets, clustering increases the NLL. For $10^{5}$ samples (or $3.1\cdot 10^{4}$ samples for the Miniboone dataset), clustering improves the NLL over any $10^{4}$ sample setting because it allows TabPFN to access a larger number of samples. However, NSF performs better than or equal to TabPFN on all datasets except the Gas dataset at $10^{5}$ samples.
@@ D.9 Unconditional density estimation on the UCI datasets
Figure D-11: Posterior predictives of TSNPE-PFN for real observations from the Allen cell type database and a simulation budget of $10^{4}$. Note that the inference is not performed directly on the action potential time series but on seven summary statistics computed from it.
@@
Figure D-12: Posterior predictives of the TSNPE baseline for real observations from the Allen cell type database and a simulation budget of $10^{4}$. Note that the inference is not performed directly on the action potential time series but on seven summary statistics computed from it.
@@
Figure D-13: Pyloric simulator. Posterior distributions for all 31 parameters estimated by TSNPE-PFN.